# Towards Autonomous Agents: Adaptive-planning, Reasoning, and Acting in Language Models

### v1

@@ V Discussion
In the failed task (task number 3), we only reflection obtained was: ”I was stuck in a loop in which I continually examined the fridge 1 instead of using a different action. I should have looked for a lettuce in the fridge 1, then taken it. I will try to execute a different action if I am stuck in a loop again.” for every trial. However, the LLM did not attempt different actions in subsequent trials and continued to examine the fridge. The actual issue in task 3 is that the fridge does not contain lettuce, but the LLM incorrectly assumes it is present and persistently attempts to retrieve it.
@@
In the sixth task (task number 6), the task could not be completed in the first trial but was successfully completed in the second trial with 38 steps. The goal of this task is to ”find two pillow and put them in sofa”. In the first trial, the agent failed to complete the task because it attempted to pick up ’pillow 2’ from ’sofa 1’; however, ’sofa 1’ only contained ’pillow 1’. Afterward, the agent continued trying to put ’pillow 2’ on ’sofa 1,’ but it failed because it did not have the pillow. Subsequently, the LLM did not output any text until the end of the trial. After this, the following adaptation was generated and appended to the input prompt for the second trial: ”I was stuck in a loop in which I continually looked for the second pillow in sofa 1. I should have looked for the second pillow in armchair 1, sidetable 1, and cabinet 1-4. I will try to execute a different action if I am stuck in a loop again.”. In the second trial, the agent found that ’pillow 2’ was on ’armchair 1,’ picked it up, placed it on the sofa, and completed the task. The trajectories of the two trials in task number 6 are shown in Fig. 6.
Figure 6: Two trajectories in task number 6 are shown in Table II. The text in each black box represents one trajectory in the ALFWorld environment [10], as detailed in the experiment in Section IV-B. In each black box, the text at the top includes the description of available receptacles and the goal instruction, as mentioned in Fig. 2. Below the goal instruction, the text following the greater-than symbol (’$>$’) represents the thoughts or actions generated by the gemma-2-9b-it LLM. Text beginning with ’$>$ think:’ indicates a thought, while all other text beginning with without ’think’ represents actions. The text beneath each thought or action is the observation from the ALFWorld environment [10]. The trajectory of the first trial in task number 6 is shown in the left black box. The LLM agent failed to complete the task because it couldn’t find the second pillow, incorrectly believing it had finished the task, and then continued generating empty strings. The adaptation the LLM agent generated is highlighted with a cyan background. In the second trial, shown in the right black box, the adaptation is appended above the description of the available receptacles. The agent successfully found the second pillow in the second trial and completed the task.
@@
In the twelfth task (task number 12), the task could not be completed in the first trial but was successfully completed in the second trial with 18 steps. The goal of this task is to ”put a cool tomato in microwave”. In the first trial, the agent failed to complete the task because it attempted to take a ’tomato’ from ’fridge 1’; however, ’fridge 1’ doesn’t contained any tomato. Afterward, the agent continued trying to take a ’tomato’ from ’fridge 1’ until the end of the trial. After this, the following adaptation was generated and appended to the input prompt for the second trial: ”I was stuck in a loop in which I continually looked for a tomato in the fridge. I should have looked for a tomato in a different environment. I will try to look for a tomato in a different environment in the next trial.”. In the second trial, the agent found that ’tomato 1’ was on ’countertop 1,’ picked it up, cooled it with the fridge, placed it in the microwave, and completed the task. The trajectories of the two trials in task number 12 are shown in Fig. 7.
Figure 7: Two trajectories in task number 12 are shown in Table II. The text in each black box represents one trajectory in the ALFWorld environment [10], as detailed in the experiment in Section IV-B. In each black box, the text at the top includes the description of available receptacles and the goal instruction, as mentioned in Fig. 2. Below the goal instruction, the text following the greater-than symbol (’$>$’) represents the thoughts or actions generated by the gemma-2-9b-it LLM. Text beginning with ’$>$ think:’ indicates a thought, while all other text beginning with without ’think’ represents actions. The text beneath each thought or action is the observation from the ALFWorld environment [10]. The trajectory of the first trial in task number 6 is shown in the left black box. The LLM agent failed to complete the task because it couldn’t find a tomato, and then continued checking the same receptacle. The adaptation the LLM agent generated is highlighted with a cyan background. In the second trial, shown in the right black box, the adaptation is appended above the description of the available receptacles. The agent successfully found a tomato in the second trial and completed the task.
VI Conclusion
@@
We present a novel in-context learning algorithm designed for a single language model to complete tasks in a text-based game by correcting its previous failures. This approach reduces the number of models used in previous work [7] from two LLMs to one LLM. Our findings indicate that the gemma-2-9b-it model achieves the highest success rate of 62% for completing tasks in the ALFWorld environment [10] using ReAct prompting [6], compared to other selected open-source language models. We show that using gemma-2-9b-it, two of the six tasks that could not be completed in one trial can be completed in the second attempt by appending the adaptation from the previous trial. Future work will involve further experimentation with different approaches to completing decision-making tasks using a single language model without Reflexion exemplars to reduce the number of tokens in the input prompt.

### v2

@@ 4.2 Experimental Result of SALA
In one of the failed tasks (task 3) from SALA, the only adaptation obtained was: "I was stuck in a loop in which I continually looked for a lettuce in the fridge. I should have looked for a lettuce in the fridge, then taken it, then put it in the countertop. I will try to execute a different action if I am stuck in a loop again." for every trial. However, the SALA agent did not attempt different actions in subsequent trials and continued to examine the fridge. The actual issue in task 3 is that the fridge does not contain lettuce, but the SALA agent incorrectly assumes it is present and persistently attempts to retrieve it. A similar issue is also found in task number 13 which has the adaptation: "I was stuck in a loop in which I continually examined fridge 1 instead of using the fridge to cool the lettuce. I should have taken the lettuce from the fridge and then put it on the countertop. I will try to execute a different action if I am stuck in a loop again." The SALA agent did not attempt different actions in subsequent trials and continued to examine the fridge.
@@
In task 6, the task could not be completed in the first trial but was successfully completed in the second trial with 38 steps. The goal of this task is to "find two pillow and put them in sofa". In the first trial, the agent failed to complete the task because it attempted to pick up ’pillow 2’ from ’sofa 1’; however, ’sofa 1’ only contained ’pillow 1’. Afterward, the agent continued trying to put ’pillow 2’ on ’sofa 1,’ but it failed because it did not have the pillow. Subsequently, the LLM did not output any text until the end of the trial. After this, the following adaptation was generated and appended to the input prompt for the second trial: "I was stuck in a loop in which I continually looked for the second pillow in sofa 1. I should have looked for the second pillow in armchair 1, sidetable 1, and cabinet 1-4. I will try to execute a different action if I am stuck in a loop again.". In the second trial, the agent found that ’pillow 2’ was on ’armchair 1,’ picked it up, placed it on the sofa, and completed the task. The trajectories of the two trials in task number 6 are shown in Figure 7 in Appendix A.1.
@@
In task 12, the task could not be completed in the first trial but was successfully completed in the second trial within 18 steps. The goal of this task is to "put a cool tomato in microwave". In the first trial, the agent failed to complete the task because it attempted to take a ’tomato’ from ’fridge 1’; however, ’fridge 1’ doesn’t contained any tomato. Afterward, the agent continued trying to take a ’tomato’ from ’fridge 1’ until the end of the trial. After this, the following adaptation was generated and appended to the input prompt for the second trial: "I was stuck in a loop in which I continually looked for a tomato in the fridge. I should have looked for a tomato in a different environment. I will try to look for a tomato in a different environment in the next trial.". In the second trial, the agent found that ’tomato 1’ was on ’countertop 1,’ picked it up, cooled it with the fridge, placed it in the microwave, and completed the task. The trajectories of the two trials in task number 12 are shown in Figure 8 in Appendix A.1.
5 Conclusion
@@
We present a novel in-context learning algorithm designed for a single language model to complete tasks in a text-based game by correcting its previous failures. This approach, by introducing self-adaptation reduces the number of models used in previous work [7] from two LLMs to one LLM, gaining in autonomy. Our findings indicate that the gemma-2-9b-it model achieves the highest success rate of 62% for completing tasks in the ALFWorld environment [10] using ReAct prompting [6], compared to other selected open-source language models. In the twelve selected tasks from the ALFWorld environment [10], the proposed SALA, utilizing the gemma-2-9b-it model, achieved a success rate of 83%, outperforming the ReAct-based agent [6] with the same model, which attained a 67% success rate. In addition, we show that using SALA with gemma-2-9b-it, two of the six tasks that could not be completed in one trial can be completed in the second attempt by appending the adaptation from the previous trial. Future work will involve further experimentation with different in-context learning algorithms to complete decision-making tasks using a single language model without Reflexion exemplars for reducing the number of tokens in the input prompt.
@@ Acknowledgments
Appendix A Appendix / supplemental material
A.1 Figures
Figure 5: Two exemplars in Reflexion [7] for the ALFWorld environment [10]. The text in each black box comprises one exemplar from Reflexion [7] designed to guide an LLM in generating the correct action to complete a task in the ALFWorld environment [10]. In each black box, the text preceding the yellow-background text represents a ReAct trajectory, as shown in Fig. 4. The text next to "STATUS: " indicates whether the task is completed. If the task is completed, the yellow-background text will read "STATUS: OK". If the task is not completed, it will read "STATUS: FAIL". The reflection text is highlighted with a cyan background.
Figure 6: The process of using a single LLM to generate cycles of thoughts, actions, and self-adaptations by fusing ReAct’s [6] and Reflexion’s exemplars [7] for solving tasks in the ALFWorld environment [10].
In Figure 7, the text in each black box represents one trajectory in the ALFWorld environment [10], as detailed in the experiment in Section 4.2. In each black box, the text at the top includes the description of available receptacles and the goal instruction, as mentioned in Fig. 3. Below the goal instruction, the text following the greater-than symbol (’$>$’) represents the thoughts or actions generated by the gemma-2-9b-it LLM. Text beginning with ’$>$ think:’ indicates a thought, while all other text beginning with without ’think’ represents actions. The text beneath each thought or action is the observation from the ALFWorld environment [10]. The trajectory of the first trial in task number 6 is shown in the left black box. The LLM agent failed to complete the task because it couldn’t find the second pillow, incorrectly believing it had finished the task, and then continued generating empty strings. The adaptation the LLM agent generated is highlighted with a cyan background. In the second trial, shown in the right black box, the adaptation is appended above the description of the available receptacles. The agent successfully found the second pillow in the second trial and completed the task.
Figure 7: Two trajectories for task 6 generated by SALA. The left trajectory represents a failure, where the agent incorrectly assumes task completion and ceases actions. The right trajectory includes an adaptation (highlighted in cyan), leading to successful task completion.
In Figure 8, the text in each black box represents one trajectory in the ALFWorld environment [10], as detailed in the experiment in Section 4.2. In each black box, the text at the top includes the description of available receptacles and the goal instruction, as mentioned in Fig. 3. Below the goal instruction, the text following the greater-than symbol (’$>$’) represents the thoughts or actions generated by the gemma-2-9b-it LLM. Text beginning with ’$>$ think:’ indicates a thought, while all other text beginning with without ’think’ represents actions. The text beneath each thought or action is the observation from the ALFWorld environment [10]. The trajectory of the first trial in task number 6 is shown in the left black box. The LLM agent failed to complete the task because it couldn’t find a tomato, and then continued checking the same receptacle. The adaptation the LLM agent generated is highlighted with a cyan background. In the second trial, shown in the right black box, the adaptation is appended above the description of the available receptacles. The agent successfully found a tomato in the second trial and completed the task.
Figure 8: Two trajectories for task 12 generated by SALA. The left trajectory shows a failure, where the agent repeats the same action with no progress until the maximum step is reached. The right trajectory includes an adaptation (in cyan), leading to successful task completion.